OpenStudy (anonymous):

Which fraction is equivalent to 5/12 ? A.2/3 B.10/18 C.20/54 D.15/36

OpenStudy (anonymous):

15/36

OpenStudy (anonymous):

i will give a metal

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with one more

OpenStudy (gabebae):

It's right

OpenStudy (anonymous):

Choose all answers that are correct. 4. (1 pt) Which fractions are equivalent to -16/24 ? Choose all answers that are correct. A.-2/3 B.-5/8 C.-32/48 D.-6/9

OpenStudy (anonymous):

A

OpenStudy (babylove18):

c

OpenStudy (texaschic101):

yep ^^ both A and C

OpenStudy (babylove18):

a,c

OpenStudy (anonymous):

ok thanks you i will tell you if is correct give me a second

OpenStudy (anonymous):

it was a,c,d

OpenStudy (anonymous):

but how did you get it

OpenStudy (babylove18):

for a its -2*8 =-16 and 3 * 8 =24 so -16/24 and c -16 *2 =-32 and 24*2=48 and d because -16/-6 =-2.6666 and 24/9=2.66666 @7thgradeink12